OpenStudy (lgbasallote):

In a certain country ½ of 5 = 3. If the same proportion holds, what is the value of 1/3 of 10 ?

OpenStudy (turingtest):

4

OpenStudy (lgbasallote):

how'd you do it? is it 5/2:3 = 10/3: x?

OpenStudy (turingtest):

\[1/2=3/5\implies1/3=2/5\]

OpenStudy (turingtest):

awww, I wanted to hear what @FoolForMath had to say :( yeah, maybe your way is more thorough, but I'm tired, and the answer's right, so ha!

OpenStudy (lgbasallote):

i wanted to hear it too..thanks anyway turing :D

OpenStudy (anonymous):

I am doing this in a single line, \[ \frac12 \times 5 = 3 \implies 2 \times \left(\frac 23 \times \left (\frac12 \times 5 \right) \right)= 2\times \left( \frac 23 \times 3\right) \implies \frac 13 \times 10 =4 \]

OpenStudy (anonymous):

\[ \frac12 \times 5 = 3 \implies 2 \times \left(\frac 23 \times \left (\frac12 \times 5 \right) \right) \] \[= 2\times \left( \frac 23 \times 3\right) \implies \frac 13 \times 10 =4 \]

OpenStudy (turingtest):

haha I beat you FFM: I did it in a single algebraic move ;)

OpenStudy (anonymous):

Good for you :P